# Primal-Dual iLQR for GPU-Accelerated Learning and Control in Legged Robots

### v2

@@ I Introduction
While the above methods have been shown to exploit the structure of the OCP, they have struggled to exploit new hardware accelerators like GPUs. Compared to CPUs, graphics cards can massively parallelize computation and are specifically designed for high-throughput linear algebra operations. Jeon et al. [cusadi] tried to bridge the gap by developing Cusadi, a tool to convert expressions written in the well-known framework CasADI [casadi] to GPU. Cusadi only translates closed-form expressions into GPU-compatible code, lacking the branching and looping capabilities necessary for most solvers. Bishop et al., [reluQP] introduced a GPU accelerated Quadratic Programming (QP) solver, Relu-QP, leveraging the Alternating Direction Method of Multipliers (ADMM) algorithm, achieving performance superior to state-of-the-art CPU-based solutions. However, its applicability to Nonlinear Programming (NLP) is constrained by the need to pre-compute offline part of the algorithm. Lee et al. [gpu_lqr_limited_parallelization] implements an iLQR controller on GPU but only partially leverages its capabilities, limiting parallelization to the line search and gradient computation. In contrast, [gpu_lqr_batched] focuses only on batching, solving multiple OCP s in parallel without exploiting state or temporal parallelism. [gpu_ddp] extends the approach of [gpu_lqr_limited_parallelization] by incorporating the temporal parallelization strategy introduced in [Mpc]. However, [Mpc] requires introducing approximations in the backward pass and performing a consensus sweep to maintain consistency. Frasch et al. [moritz_parallel], proposed a method to iteratively solve linear quadratic control exploiting parallel computation, while Wright et al. [partitioned_dynamics] showed how to partition dynamic programming for parallel computation. However, none of the mentioned methods achieved at the same time logarithmic time complexity and an exact formulation. On the other hand, Särkkä et al. [TemporalParallelizationLQR] showed how parallel associative scan operations can be utilized to improve the computational complexity of DDP. When parallelized on GPU, associative scans can compute the optimal control policy in $\mathcal{O}(\log^{2}(n)\log{N}+\log^{2}(m))$ instead of the $\mathcal{O}(N(n+m)^{3})$ of the classical Riccati Recursion found in DDP, where $n$ and $m$ are the dimensions of the system state and control input, and $N$ is the prediction horizon length.
@@
In this work, we designed an SQP method that uses an associative scan-based LQR solver to solve the primal-dual Karush-Kuhn-Tucker (KKT) system efficiently. The solver avoids the offline precalculation necessary for Relu-QP [reluQP]. Our method benefits from a multiple shooting implementation, instead of the single shooting approach presented in [TemporalParallelizationLQR]. Our algorithm fully exploits the parallelization capabilities of GPUs both in the temporal and the state space dimensions, while maintaining, in contrast to [gpu_ddp], an exact backward pass. We tailored the formulation for the deployment as a receding horizon controller for legged robot locomotion and analyzed its impact on learning and control.
@@
Control:  At first, the reduction in the computational complexity of the solver may appear to have limited significance in a practical application. Some of the already described methods, like OCS2 [ocs2] and crocoddyl [crocoddyl], have been successfully used for whole-body MPC even for experiments, as shown in [inverse_dynamics_mpc] and [Perceptive_based_MPC]. To ensure that such an MPC formulation can run online on real hardware, the prediction horizon length and the number of robots that can be controlled simultaneously are constrained by the time complexity of the underlying algorithm. In particular, for solvers like OCS2, acados, and crocoddyl, the complexity scales linearly with the horizon length and cubically with the state dimension. In contrast, our approach scales with the square-log of the horizon length and linearly with the state dimension. This improved scaling enables the use of more nodes at the same update rate, allowing for finer integration steps for a better approximation of the robot dynamic or, more critically, the inclusion of a Moving Horizon Estimator (MHE), such as those presented by [Invariant_moother] or [inertial_estimationlocalization], within the same optimization loop, opening the pace for ”end-to-end” MPC s capable of adapting and reacting to system changes directly using the sensor data. The reduction in complexity with respect to the state dimension enables better scalability across different robot morphologies. While other methods, such as the one presented by [AcceleratingMPC], use an approximate approach in the form of consensus ADMM to parallelize over the system state, our implementation leverages the inherent parallelism of matrix operations provided by the GPU. State space scaling becomes crucial when a centralized controller for collaborative tasks is considered. As demonstrated by [CentralizedMPC] and [LayeredMPC], centralized controllers can handle complex tasks involving multiple collaborating agents. However, while these approaches achieve notable performances, the complexity of the task and the number of agents involved are constrained by their capability to solve the increasingly large OCP s at a reasonable control frequency. Our framework overcomes such barriers, as shown in Sec. IV-C.
@@ I-A Contribution
  • a novel GPU-accelerated MPC for legged robots that achieves logarithmic scaling in computation complexity on the horizon length, linear scaling in the state space and control vector dimensions, and can easily be parallelized for use with data-driven approaches.
  • a detailed analysis of the performance benefits and limitations of the proposed algorithm with respect to state-of-the-art solvers.
  • an open-source code repository for the rapid prototyping of MPC for legged robots and loco-manipulation. Written in JAX, the code provides tools for controlling legged robots and also supports large-scale parallelization.
@@ II-A Optimal Control Problem
We start describing the OCP at the base of the MPC formulation we are presenting. In a OCP, we transcribe not only the task we want to be performed but also the physical limits the system needs to respect. In this paper, we consider the OCP in the following form:
@@
$s.t.\quad$ $\boldsymbol{x}_{i+1}=\boldsymbol{h}(\boldsymbol{x}_{i},\boldsymbol{u}_{i})$ (1b)
@@ II-B Sequential Quadratic Programming
We start defining the Lagrangian relative to the problem in (1) as:
$\mathcal{L}(\boldsymbol{x},\boldsymbol{u},\boldsymbol{\lambda})=$ $\boldsymbol{l}_{N+1}(\boldsymbol{x}_{N+1})+\sum_{i=0}^{N}{\boldsymbol{l}_{i}(\boldsymbol{x}_{i},\boldsymbol{u}_{i})}+$ (2)
$\boldsymbol{{\color[rgb]{0,0,0}\lambda_{i}}}^{T}(\boldsymbol{x}_{i+1}-\boldsymbol{h}(\boldsymbol{x}_{i},\boldsymbol{u}_{i}))$
where $\boldsymbol{\lambda}_{i}$ for $i=1\,\ldots\,N+1$ are the Lagrangian multipliers related to the dynamics. In an SQP method, we iteratively solve the QP derived from the quadratic approximation of the Lagrangian and the linearization of the constraints at the current guess, $\boldsymbol{x}^{k}$, $\boldsymbol{u}^{k}$, where $k$ represents the last iterate. The QP is then written as:
@@
$\boldsymbol{\delta x}_{0}=0$
$\boldsymbol{\delta x}_{i+1}=\boldsymbol{A}_{i}\boldsymbol{\delta x}_{i}+\boldsymbol{B}_{i}\boldsymbol{\delta u}_{i}+\boldsymbol{b}_{i}~~\text{for}\quad i={\color[rgb]{0,0,0}0},\dots,N$
where the search directions $\boldsymbol{\delta x}$ and $\boldsymbol{\delta u}$ are defined as:
@@
The linearized dynamics is defined as:
$\boldsymbol{A}_{i}=\frac{\partial\boldsymbol{h}_{i}}{\partial\boldsymbol{x}_{i}},\,\boldsymbol{B}_{i}=\frac{\partial\boldsymbol{h}_{i}}{\partial\boldsymbol{u}_{i}},\,\boldsymbol{b}_{i}=\boldsymbol{h}_{i}-\boldsymbol{x}^{k}_{i+1}$
The linear terms in the objective function are defined as:
@@
Finally, the Quadratic terms are defined as:
$\boldsymbol{P}_{N+1}=\nabla^{2}\mathcal{L}_{N+1},\quad\begin{bmatrix}\boldsymbol{Q}_{i}&\boldsymbol{S}_{i}^{\top}\\ \boldsymbol{S}_{i}&\boldsymbol{R}_{i}\end{bmatrix}=\nabla^{2}\mathcal{L}_{i}$
where we removed the dependency of $\boldsymbol{x}^{k}_{i}$ and $\boldsymbol{u}^{k}_{i}$ from the dynamics $\boldsymbol{h}$ and the cost function $\boldsymbol{l}$.
@@ II-C Primal & dual problem solution
As shown by [fastGeneration], such QP can be efficiently solved by exploiting the problem structure. By applying the Riccati Recursion, starting from the $N$ last node in the horizon and going backward in time we obtain:
@@
$\boldsymbol{H}_{i}$ $=\boldsymbol{B}_{i}^{\top}\boldsymbol{P}_{i+1}\boldsymbol{A}_{i},$
@@
$\boldsymbol{P}_{i}$ $=\boldsymbol{Q}_{i}+\boldsymbol{A}_{i}^{\top}\boldsymbol{P}_{i+1}\boldsymbol{A}_{i}+(\boldsymbol{S}_{i}^{\top}+\boldsymbol{A}_{i}\boldsymbol{P}_{i+1}\boldsymbol{B}_{i})\boldsymbol{K}_{i},$
@@
where $\boldsymbol{K}_{i}$ and $\boldsymbol{k}_{i}$ are, respectively the feedback and feed-forward term of the optimal control policy. We can now retrieve the control and state vector applying recursively the forward pass starting from the initial condition as follows:
$\boldsymbol{\delta u}_{i}=\boldsymbol{K}_{i}\boldsymbol{\delta x_{i}}+\boldsymbol{k}_{i}\quad\boldsymbol{\delta x}_{i+1}=\boldsymbol{A}_{i}\boldsymbol{\delta x}_{i}+\boldsymbol{B}\boldsymbol{\delta u}_{i}+\boldsymbol{b}_{i}$ (5)
Once retrieved the values of $\boldsymbol{P}_{i}$ and $\boldsymbol{p}_{i}$ for $i=0,\dots,N+1$ we can compute the update of the Lagrangian multiplier $\boldsymbol{\delta\lambda}$ in parallel. This is achieved by applying, at each node, the following update:
$\boldsymbol{\delta\lambda}_{i}=\boldsymbol{P}_{i}\boldsymbol{\delta x}_{i}+\boldsymbol{p}_{i}$ (6)
@@ II-D Parallel Associative Scan
in $\mathcal{O}(\log{N})$ as shown by [scan]. This is possible since $\otimes$ is an associative operation, allowing the computation to be reorganized into smaller interdependent subproblems. To make use of this property, we first define the Conditional Value Function (CVF) $V_{i\rightarrow{}j}(\boldsymbol{x}_{i},\boldsymbol{x}_{j})$ as the minimal cost related to the optimal trajectory that goes from state $\boldsymbol{x}_{i}$ at node $i$ in the horizon to state $\boldsymbol{x}_{j}$ at node $j$, with $i<j$. Given the connection between the value function of (II-B) and its CVF, $V_{i}(\boldsymbol{x}_{i})=V_{i\rightarrow N+1}(\boldsymbol{x}_{i},\boldsymbol{x}_{N+1})$, we will use the scan to evaluate $V_{i}(\boldsymbol{x}_{i})$ for $i=0,\dots N+1$. We begin by initializing the elements $a_{i}$ with the CVF $V_{i\rightarrow{i+1}}$ defined between two adjacent states $\boldsymbol{x}_{i}$ and $\boldsymbol{x}_{i+1}$. Next, we define the operator $\otimes$ for dynamic programming as:
@@
$V_{i\rightarrow{}j}(\boldsymbol{x}_{i},\boldsymbol{x}_{j})=\max\limits_{\boldsymbol{\eta}}$ $\frac{1}{2}\boldsymbol{x}_{i}^{\top}\boldsymbol{\tilde{P}}_{i,j}\boldsymbol{x}_{i}+\boldsymbol{\tilde{p}}_{i,j}^{\top}\boldsymbol{x}_{i}-\frac{1}{2}\boldsymbol{{\eta}}^{\top}\boldsymbol{\tilde{C}}_{i,j}\boldsymbol{{\eta}}+$ (9)
@@
where the tilde $\tilde{\cdot}$ is used to distinguish the values related to the conditional value function from the one in (4), while the subscripts are used to specify the initial and final state considered. Can be proved that given $V_{i\rightarrow{k}}$ and $V_{k\rightarrow{j}}$ that are in the same form of (9), $V_{i\rightarrow{j}}$ will still be in the form of (9) and is characterized by such combination rule:
@@
We now initialize the values $a_{i}$ of a reverse associative scan, using the conditional value function $V_{i\rightarrow{i+1}}$ characterized by:
$\boldsymbol{\tilde{A}}_{i,i+1}$ $=\boldsymbol{A_{i}}-\boldsymbol{B_{i}}\boldsymbol{R}_{i}^{-1}S_{i}^{\top},$ $\boldsymbol{\tilde{P}}_{i,i+1}$ $=\boldsymbol{Q_{i}}-\boldsymbol{S_{i}}\boldsymbol{R}_{i}^{-1}\boldsymbol{S}_{i}^{\top}$ (11)
$\boldsymbol{\tilde{C}}_{i,i+1}$ $=\boldsymbol{B_{i}}\boldsymbol{R_{i}}^{-1}\boldsymbol{B}_{i}^{\top},$ $\boldsymbol{\tilde{p}}_{i,i+1}$ $=\boldsymbol{q_{i}}-\boldsymbol{S_{i}}\boldsymbol{R}_{i}^{-1}r_{i}$
@@
while $V_{N,N+1}$ by :
$\boldsymbol{\tilde{P}}_{N,N+1}$ $=\boldsymbol{Q}_{N+1},$ $\boldsymbol{\tilde{A}}_{N,N+1}$ $=0,$ $\boldsymbol{\tilde{b}}_{N,N+1}=0$ (12)
$\boldsymbol{\tilde{p}}_{N,N+1}$ $=\boldsymbol{q}_{N+1}$ $\boldsymbol{\tilde{C}}_{N,N+1}$ $=0$
@@
From the solution of the parallel associative scan using the combination rule in (10), we obtain $V_{i\rightarrow N+1}(\boldsymbol{x}_{i},\boldsymbol{x}_{N+1})$ for $i=0,\dots,N+1$. Finally, recalling that $V_{i\rightarrow{N+1}}(\boldsymbol{x}_{i},\boldsymbol{x}_{N+1})$ is equal to the value function $V_{i}(x_{i})$ of the problem in (II-B) and that the value function is also in the same form of (9) we get that $\boldsymbol{P}_{i}=\boldsymbol{\tilde{P}}_{i,N+1}$ and $\boldsymbol{\tilde{p}}_{i}=\boldsymbol{p}_{i,N+1}$. This means that we obtained all $\boldsymbol{P}_{i}$ and $\boldsymbol{p}_{i}$ for $i=0,\,\dots\,,N+1$ that can then be used in (II-C) to calculate $\boldsymbol{K}$ and $\boldsymbol{k}$ in on pass in parallel. Thanks to a GPU-accelerated scan implementation, we can obtain the control policy over the entire horizon with an overall computational complexity of $\mathcal{O}(\log^{2}(n)\log{N}+\log^{2}(m))$. This follows from the fact that both computing and initializing the scan involve a matrix inversion, which can be performed in parallel with complexity $\mathcal{O}(\log^{2}(n))$ [parallel_matrix_inv]. The parallel associative scan can also be used to retrieve the optimal trajectory in $\mathcal{O}(\log{n}\log{N}+\log{m})$. We start plugging into the linearized dynamics the optimal control law we evaluated, obtaining:
$\boldsymbol{\delta x}_{i+1}=\boldsymbol{\bar{A}}_{i}\boldsymbol{\delta x}_{i}+\boldsymbol{\bar{b}}_{i}$ (13)
where: $\boldsymbol{\bar{A}}_{i}=\boldsymbol{A}_{i}+\boldsymbol{B}_{i}\boldsymbol{K}_{i}$ and $\boldsymbol{\bar{b}}_{i}=\boldsymbol{B}_{i}\boldsymbol{k}_{i}+\boldsymbol{b}_{i}$ We can define the conditional optimal trajectory as follows:
$\boldsymbol{\bar{h}}_{i\rightarrow{j}}(\boldsymbol{\delta x}_{i},\boldsymbol{\delta x}_{j})=\boldsymbol{\bar{A}}_{i,j}\boldsymbol{\delta x}_{i}+\boldsymbol{\bar{b}}_{i,j}$ (14)
@@
[TemporalParallelizationLQR] showed that from the combination of two conditional optimal trajectory $\boldsymbol{h}_{i\rightarrow{k}}$ and $\boldsymbol{h}_{k\rightarrow{j}}$ we obtain $\boldsymbol{h}_{i\rightarrow{j}}$ that can be written in the same form as (14) with:
$\boldsymbol{\bar{A}}_{i,j}=\boldsymbol{\bar{A}}_{i,k}\boldsymbol{\bar{A}}_{k,j}\quad\boldsymbol{\bar{b}}_{i,j}=\boldsymbol{\bar{A}}_{k,j}\boldsymbol{\bar{b}}_{i,k}+\boldsymbol{\bar{b}}_{k,j}$ (15)
We can initialize the element of parallel scan $a_{i}=\boldsymbol{h}_{i\rightarrow{i+1}}(\boldsymbol{x}_{i})$ as:
$\boldsymbol{\bar{A}}_{i,i+1}=\boldsymbol{\bar{A}}_{i}\quad\boldsymbol{\bar{b}}_{i,i+1}=\boldsymbol{\bar{b}}_{i}\\$ (16)

### v3

@@ I Introduction
While the above methods have been shown to exploit the structure of the OCP, they have struggled to exploit new hardware accelerators like GPUs. Compared to CPUs, graphics cards can massively parallelize computation and are specifically designed for high-throughput linear algebra operations. Jeon et al. [cusadi] tried to bridge the gap by developing Cusadi, a tool to convert expressions written in the well-known framework CasADI [casadi] to GPU. Cusadi only translates closed-form expressions into GPU-compatible code, lacking the branching and looping capabilities necessary for most solvers. Bishop et al., [reluQP] introduced a GPU accelerated Quadratic Programming (QP) solver, Relu-QP, leveraging the Alternating Direction Method of Multipliers (ADMM) algorithm, achieving performance superior to state-of-the-art CPU-based solutions. However, its applicability to Nonlinear Programming (NLP) is constrained by the need to pre-compute part of the algorithm offline. Lee et al. [gpu_lqr_limited_parallelization] implements an iLQR controller on GPU but only partially leverages its capabilities, limiting parallelization to the line search and gradient computation. In contrast, [gpu_lqr_batched] focuses only on batching, solving multiple OCP s in parallel without exploiting state or temporal parallelism. [gpu_ddp] extends the approach of [gpu_lqr_limited_parallelization] by incorporating the temporal parallelization strategy introduced in [Mpc]. However, [Mpc] requires introducing approximations in the backward pass and performing a consensus sweep to maintain consistency. Frasch et al. [moritz_parallel], proposed a method to iteratively solve linear quadratic control exploiting parallel computation, while Wright et al. [partitioned_dynamics] showed how to partition dynamic programming for parallel computation. However, none of the mentioned methods achieved at the same time logarithmic time complexity and an exact formulation. On the other hand, Särkkä et al. [TemporalParallelizationLQR] showed how parallel associative scan operations can be utilized to improve the computational complexity of LQR. When parallelized on a GPU, associative scans can compute the optimal control policy in $\mathcal{O}(\log^{2}(n)\log{N}+\log^{2}(m))$ instead of the $\mathcal{O}(N(n+m)^{3})$ of the classical Riccati Recursion, where $n$ and $m$ are the dimensions of the system state and control input, and $N$ is the prediction horizon length.
@@
In this work, we designed an SQP method that uses an associative scan-based LQR solver to solve the primal-dual Karush-Kuhn-Tucker (KKT) system efficiently. The solver avoids the offline precalculation necessary for Relu-QP [reluQP]. Our method benefits from a multiple shooting implementation, instead of the single shooting approach presented in [gpu_ddp]. Our algorithm fully exploits the parallelization capabilities of GPUs in the temporal, state, and control dimensions, while maintaining, in contrast to [gpu_ddp], an exact backward pass. We tailored the formulation for the deployment as a receding horizon controller for legged robot locomotion and analyzed its impact on learning and control.
@@
Control:  At first, the reduction in the computational complexity of the solver may appear to have limited significance in a practical application. Some of the already described methods, like OCS2 [ocs2] and crocoddyl [crocoddyl], have been successfully used for whole-body MPC even for experiments, as shown in [inverse_dynamics_mpc] and [Perceptive_based_MPC]. To ensure that such an MPC formulation can run online on real hardware, the prediction horizon length and the number of robots that can be controlled simultaneously are constrained by the time complexity of the underlying algorithm. In particular, for solvers like OCS2, acados, and crocoddyl, the complexity scales linearly with the horizon length and cubically with the state dimension. In contrast, our approach scales with the square-log of the horizon length and logarithmically with the state and control dimensions. This improved scaling enables the use of more nodes at the same update rate, allowing for finer integration steps for a better approximation of the robot dynamic or, more critically, the inclusion of a Moving Horizon Estimator (MHE), such as those presented by [Invariant_moother] or [inertial_estimationlocalization], within the same optimization loop, opening the pace for ”end-to-end” MPC s capable of adapting and reacting to system changes directly using the sensor data. The reduction in complexity with respect to the state dimension enables better scalability across different robot morphologies. While other methods, such as the one presented by [AcceleratingMPC], use an approximate approach in the form of consensus ADMM to parallelize over the system state, our implementation leverages the inherent parallelism of matrix operations provided by the GPU. State space scaling becomes crucial when a centralized controller for collaborative tasks is considered. As demonstrated by [CentralizedMPC] and [LayeredMPC], centralized controllers can handle complex tasks involving multiple collaborating agents. However, while these approaches achieve notable performances, the complexity of the task and the number of agents involved are constrained by their capability to solve the increasingly large OCP s at a reasonable control frequency. Our framework overcomes such barriers, as shown in Sec. IV-C.
@@ I-A Contribution
  • a novel GPU-accelerated MPC for legged robots that achieves logarithmic scaling in computation complexity on the horizon length, squared-logarithmic scaling in the state and control dimensions, and can easily be parallelized for use with data-driven approaches;
  • a detailed analysis of the performance benefits and limitations of the proposed algorithm with respect to state-of-the-art solvers;
  • an open-source code repository for the rapid prototyping of MPC for legged robots and loco-manipulation, written in JAX; the code provides tools for controlling legged robots and also supports large-scale parallelization.
@@ II-A Optimal Control Problem
We start describing the OCP at the base of the MPC formulation we are presenting. In a OCP, we transcribe not only the task we want to be performed but also the physical limits the system needs to respect. In this paper, we consider OCP s of the following form:
@@
$s.t.\quad$ $\boldsymbol{x}_{i+1}=\boldsymbol{h}_{i}(\boldsymbol{x}_{i},\boldsymbol{u}_{i})$ (1b)
@@ II-B Sequential Quadratic Programming
We start by defining the Lagrangian of problem in (1) as
$\mathcal{L}(\boldsymbol{x},\boldsymbol{u},\boldsymbol{\lambda})=$ $\boldsymbol{l}_{N+1}(\boldsymbol{x}_{N+1})+\boldsymbol{\lambda}_{0}^{T}(\boldsymbol{\hat{x}}_{0}-\boldsymbol{x}_{0})+$ (2)
$\sum_{i=0}^{N}{\boldsymbol{l}_{i}(\boldsymbol{x}_{i},\boldsymbol{u}_{i})}+\boldsymbol{\lambda}_{i+1}^{T}(\boldsymbol{h}_{i}(\boldsymbol{x}_{i},\boldsymbol{u}_{i})-\boldsymbol{x}_{i+1}),$
where $\boldsymbol{\lambda}_{i}$ for $i=1\,\ldots\,N+1$ are the Lagrangian multipliers associated with the dynamics constraints (also known as costates). In an SQP method, we iteratively solve the QP derived from the quadratic approximation of the Lagrangian and the linearization of the constraints at the current guess, $\boldsymbol{x}^{k}$, $\boldsymbol{u}^{k}$, where $k$ represents the last iterate. The QP is then written as
@@
$\boldsymbol{\delta x}_{0}=\boldsymbol{\hat{x}}_{0}-\boldsymbol{x}_{0}^{k}$
$\boldsymbol{\delta x}_{i+1}=\boldsymbol{A}_{i}\boldsymbol{\delta x}_{i}+\boldsymbol{B}_{i}\boldsymbol{\delta u}_{i}+\boldsymbol{b}_{i}~~\text{for}\quad i={\color[rgb]{0,0,0}0},\dots,N,$
where the search directions $\boldsymbol{\delta x}$ and $\boldsymbol{\delta u}$ are defined as
@@
The linearized dynamics is defined as
$\boldsymbol{A}_{i}=\frac{\partial\boldsymbol{h}_{i}}{\partial\boldsymbol{x}_{i}},\,\boldsymbol{B}_{i}=\frac{\partial\boldsymbol{h}_{i}}{\partial\boldsymbol{u}_{i}},\,\boldsymbol{b}_{i}=\boldsymbol{h}_{i}-\boldsymbol{x}^{k}_{i+1}.$
The linear terms in the objective function are defined as
@@
Finally, the Quadratic terms are defined as
$\boldsymbol{P}_{N+1}=\nabla^{2}\mathcal{L}_{N+1},\quad\begin{bmatrix}\boldsymbol{Q}_{i}&\boldsymbol{S}_{i}^{\top}\\ \boldsymbol{S}_{i}&\boldsymbol{R}_{i}\end{bmatrix}=\nabla^{2}\mathcal{L}_{i},$
where we removed the dependency of $\boldsymbol{x}^{k}_{i}$ and $\boldsymbol{u}^{k}_{i}$ from the dynamics $\boldsymbol{h}_{i}$ and the cost function $\boldsymbol{l}_{i}$.
@@ II-C Primal & dual problem solution
As shown by [fastGeneration], such QPs can be efficiently solved by exploiting the problem structure. By applying the Riccati Recursion, starting from the last (i.e. with index $N$) node in the horizon and going backward in time, we obtain:
@@
$\boldsymbol{H}_{i}$ $=\boldsymbol{S}_{i}+\boldsymbol{B}_{i}^{\top}\boldsymbol{P}_{i+1}\boldsymbol{A}_{i},$
@@
$\boldsymbol{P}_{i}$ $=\boldsymbol{Q}_{i}+\boldsymbol{A}_{i}^{\top}\boldsymbol{P}_{i+1}\boldsymbol{A}_{i}+\boldsymbol{K}_{i}^{\top}\boldsymbol{H}_{i},$
@@
where $\boldsymbol{K}_{i}$ and $\boldsymbol{k}_{i}$ are, respectively, the feedback and feed-forward term of the optimal control policy. We can now retrieve the control and state vectors by recursively applying the forward pass starting from the initial condition as follows:
$\boldsymbol{\delta u}_{i}=\boldsymbol{K}_{i}\boldsymbol{\delta x_{i}}+\boldsymbol{k}_{i},\quad\boldsymbol{\delta x}_{i+1}=\boldsymbol{A}_{i}\boldsymbol{\delta x}_{i}+\boldsymbol{B}_{i}\boldsymbol{\delta u}_{i}+\boldsymbol{b}_{i}.$ (5)
Once we retrieve the values of $\boldsymbol{P}_{i}$ and $\boldsymbol{p}_{i}$ for $i=0,\dots,N+1$, we can compute the update of the Lagrangian multiplier $\boldsymbol{\delta\lambda}$ in parallel. This is achieved by applying, at each node, the following update:
$\boldsymbol{\delta\lambda}_{i}=\boldsymbol{P}_{i}\boldsymbol{\delta x}_{i}+\boldsymbol{p}_{i}.$ (6)
@@ II-D Parallel Associative Scan
in $\mathcal{O}(\log{N})$ as shown by [scan]. This is possible since $\otimes$ is an associative operation, allowing the computation to be reorganized into smaller interdependent subproblems. To make use of this property, we first define the Conditional Value Function (CVF) $V_{i\rightarrow{}j}(\boldsymbol{x}_{i},\boldsymbol{x}_{j})$ as the minimal cost related to the optimal trajectory that goes from state $\boldsymbol{x}_{i}$ at node $i$ in the horizon to state $\boldsymbol{x}_{j}$ at node $j$, with $i<j$. Given the connection between the value function of (II-B) and its CVF, $V_{i}(\boldsymbol{x}_{i})=\max_{\boldsymbol{x}_{N+1}}V_{i\rightarrow N+1}(\boldsymbol{x}_{i},\boldsymbol{x}_{N+1})$, we will use the scan to evaluate $V_{i}(\boldsymbol{x}_{i})$ for $i=0,\dots N+1$. We begin by initializing the elements $a_{i}$ with the CVF $V_{i\rightarrow{i+1}}$ defined between two adjacent states $\boldsymbol{x}_{i}$ and $\boldsymbol{x}_{i+1}$. Next, we define the operator $\otimes$ for dynamic programming as:
@@
$V_{i\rightarrow{}j}(\boldsymbol{x}_{i},\boldsymbol{x}_{j})=\max\limits_{\boldsymbol{\eta}}$ $\frac{1}{2}\boldsymbol{x}_{i}^{\top}\boldsymbol{\tilde{P}}_{i,j}\boldsymbol{x}_{i}+\boldsymbol{\tilde{p}}_{i,j}^{\top}\boldsymbol{x}_{i}-\frac{1}{2}\boldsymbol{{\eta}}^{\top}\boldsymbol{\tilde{C}}_{i,j}\boldsymbol{{\eta}}$ (9)
@@
where the tilde $\tilde{\cdot}$ is used to distinguish the values related to the conditional value function from the one in (4), while the subscripts are used to specify the initial and final state considered. It can be proved that, given $V_{i\rightarrow{k}}$ and $V_{k\rightarrow{j}}$ in the form of (9), $V_{i\rightarrow{j}}$ will still have the form of (9) and is characterized by the combination rule
@@
We now initialize the values $a_{i}$ of a reverse associative scan, using the conditional value function $V_{i\rightarrow{i+1}}$ characterized by
$\boldsymbol{\tilde{A}}_{i,i+1}$ $=\boldsymbol{A}_{i}-\boldsymbol{B}_{i}\boldsymbol{R}_{i}^{-1}\boldsymbol{S}_{i},$ $\boldsymbol{\tilde{P}}_{i,i+1}$ $=\boldsymbol{Q}_{i}-\boldsymbol{S}_{i}^{\top}\boldsymbol{R}_{i}^{-1}\boldsymbol{S}_{i},$ (11)
$\boldsymbol{\tilde{C}}_{i,i+1}$ $=\boldsymbol{B}_{i}\boldsymbol{R_{i}}^{-1}\boldsymbol{B}_{i}^{\top},$ $\boldsymbol{\tilde{p}}_{i,i+1}$ $=\boldsymbol{q}_{i}-\boldsymbol{S}_{i}^{\top}\boldsymbol{R}_{i}^{-1}r_{i},$
@@
and, in the case of $V_{N,N+1}$, by
$\boldsymbol{\tilde{P}}_{N,N+1}$ $=\boldsymbol{Q}_{N+1},$ $\boldsymbol{\tilde{A}}_{N,N+1}$ $=0,$ $\boldsymbol{\tilde{b}}_{N,N+1}=0,$ (12)
$\boldsymbol{\tilde{p}}_{N,N+1}$ $=\boldsymbol{q}_{N+1},$ $\boldsymbol{\tilde{C}}_{N,N+1}$ $=0.$
@@
From the solution of the parallel associative scan using the combination rule in (10), we obtain $V_{i\rightarrow N+1}(\boldsymbol{x}_{i},\boldsymbol{x}_{N+1})$ for $i=0,\dots,N+1$. Finally, recalling that $\max_{\boldsymbol{x}_{N+1}}V_{i\rightarrow{N+1}}(\boldsymbol{x}_{i},\boldsymbol{x}_{N+1})$ is equal to the value function $V_{i}(x_{i})$ of the problem in (II-B) and that the value function is also in the same form of (9), we get that $\boldsymbol{P}_{i}=\boldsymbol{\tilde{P}}_{i,N+1}$ and $\boldsymbol{\tilde{p}}_{i}=\boldsymbol{p}_{i,N+1}$. This means that we obtained all $\boldsymbol{P}_{i}$ and $\boldsymbol{p}_{i}$ for $i=0,\,\dots\,,N+1$ that can then be used in (II-C) to calculate $\boldsymbol{K}$ and $\boldsymbol{k}$ in on pass in parallel. Thanks to a GPU-accelerated scan implementation, we can obtain the control policy over the entire horizon with an overall computational complexity of $\mathcal{O}(\log^{2}(n)\log{N}+\log^{2}(m))$. This follows from the fact that both computing and initializing the scan involve a matrix inversion, which can be performed in parallel with complexity $\mathcal{O}(\log^{2}(n))$ [parallel_matrix_inv]. The parallel associative scan can also be used to retrieve the optimal trajectory in $\mathcal{O}(\log{n}\log{N}+\log{m})$. We start by plugging into the linearized dynamics the optimal control law we evaluated, obtaining
$\boldsymbol{\delta x}_{i+1}=\boldsymbol{\bar{A}}_{i}\boldsymbol{\delta x}_{i}+\boldsymbol{\bar{b}}_{i},$ (13)
where $\boldsymbol{\bar{A}}_{i}=\boldsymbol{A}_{i}+\boldsymbol{B}_{i}\boldsymbol{K}_{i}$ and $\boldsymbol{\bar{b}}_{i}=\boldsymbol{B}_{i}\boldsymbol{k}_{i}+\boldsymbol{b}_{i}$. We can define the conditional optimal trajectory as follows:
$\boldsymbol{\bar{h}}_{i\rightarrow{j}}(\boldsymbol{\delta x}_{i},\boldsymbol{\delta x}_{j})=\boldsymbol{\bar{A}}_{i,j}\boldsymbol{\delta x}_{i}+\boldsymbol{\bar{b}}_{i,j}.$ (14)
@@
[TemporalParallelizationLQR] showed that from the combination of two conditional optimal trajectory $\boldsymbol{h}_{i\rightarrow{k}}$ and $\boldsymbol{h}_{k\rightarrow{j}}$ we obtain $\boldsymbol{h}_{i\rightarrow{j}}$ that can be written in the same form as (14) with
$\boldsymbol{\bar{A}}_{i,j}=\boldsymbol{\bar{A}}_{i,k}\boldsymbol{\bar{A}}_{k,j}\quad\boldsymbol{\bar{b}}_{i,j}=\boldsymbol{\bar{A}}_{k,j}\boldsymbol{\bar{b}}_{i,k}+\boldsymbol{\bar{b}}_{k,j}.$ (15)
We can initialize the element of parallel scan $a_{i}=\boldsymbol{h}_{i\rightarrow{i+1}}(\boldsymbol{x}_{i})$ as
$\boldsymbol{\bar{A}}_{i,i+1}=\boldsymbol{\bar{A}}_{i},\quad\boldsymbol{\bar{b}}_{i,i+1}=\boldsymbol{\bar{b}}_{i},\\$ (16)